OpenStudy (anonymous):

A company makes jewelry. It costs $6,400 to make 10 pieces of jewelry and $11,200 to make 20 pieces of jewelry. Which equation models the cost, C(x), as a linear function of the number of pieces of jewelry, x?

OpenStudy (anonymous):

C(x) = 480x + 1600 C(x) = 1600x + 480 C(x) = 480x − 1600 C(x) = 1600x − 480

OpenStudy (anonymous):